for each of the following, state the condition on a and k such that the parabola y=a(x-h)^2+k has the given property.

a) The parabola intersects the x-axis at two distinct points
b)The parabola intersects the x-axis at one point.
c)The parabola does not intersect the x-axis.
1
Expert's answer
2020-12-11T01:42:53-0500

Consider the parabola "y=a(x-h)^2+k"


a) The parabola intersect the "x" -axis on setting "y=0"


So,


"a(x-h)^2+k=0"


"a(x-h)^2=-k"


"(x-h)^2=\frac{-k}{a}"


"x=h\pm\sqrt{-\frac{k}{a}}"


For real values of "x" , the term under square root sign must be positive.


Therefore, the condition for two distinct points are:

i) For "a>0" , the value of "k" should be negative, that is "k<0"

ii) For "k>0", the value of "a" should be negative, that is "a<0"


b) For "k=0" , the "x" -intercept is,

"x=h\pm\sqrt{-\frac{0}{a}}"


"=h\pm0"


"=h"


Therefore, the parabola intersect the "x" -axis at one point when "k=0" irrespective of the value of "a" .


c) Here,

"x=h\pm\sqrt{-\frac{k}{a}}"


The "x" value will be undefined if "\frac{k}{a}>0"


Therefore, the condition for the parabola not intersecting the "x" -axis are:


i) For "a>0" , the value of "k" should be positive, that is "k>0"

ii) For "a<0", the value of "k" should be negative, that is "k<0"